Taking a BigBite Out of Short Range Correlations

Peter Monaghan^{1}, Ran Shneor^{2} and Ramesh Subedi^{3} for the E01-015 Collaboration

1 Massachusetts Institute of Technology, Cambridge, MA 02139, USA 2 Tel Aviv University, Tel Aviv, 69978, Israel 3 Kent State University, Kent, Ohio, 44242, USA

Abstract

A short range correlation between a pair of nucleons occurs when their wavefunctions strongly overlap. The experiment E01-015 took data from January through April this year in Hall A at JLab, searching for short range correlations in carbon. In the impulse approximation, a virtual photon is absorbed by one nucleon of the pair, knocking it out of the nucleus. To conserve momentum, the partner nucleon recoils with its initial momentum and is also ejected from the nucleus. Although short range correlations can manifest between neutron-neutron, proton-neutron and proton-proton pairs, we focus only on the proton-proton channel here. A triple coincidence measurement using both high resolution spectrometers in Hall A and a third, large acceptance spectrometer called BigBite was made. The scattered electron was detected in the left HRS, the struck proton was detected in the right HRS and the recoil partner nucleon was detected in BigBite. The kinematics were chosen to minimize the effects of final state interactions, meson exchange currents and isobar configurations, which could dominate the cross-section over the short range correlation effects. Some preliminary data establishing triple coincidence events detected by BigBite is shown.